# Ultraviolet spectroscopy reveals a hot and luminous companion to the Be star + black hole candidate MWC 656

### v1

@@ 8.2 Caveats of the composite model
A notable feature of the UV data is an apparent Doppler shift in the high-ionisation wind lines. In three of the four epochs, the lines show no significant shift, but at the red edge of the absorption component, clearly seen, for example, in C iv (Fig. 4), appears redshifted by 80 km s-1. Measurement of RVs in wind lines is inherently difficult due to their complex and variable profiles, and the situation could be further complicated here by interstellar absorption overlapping with the wind features. Nevertheless, the observed shift cannot easily be explained by the orbital motion of either star: The amplitude is too small for the companion, the sign is opposite to the expected motion at this phase, and the shift is too large to be attributed to the Be star. This suggests that the high-ionisation line-forming region does not simply trace the orbital motion of either component. Redshifted absorption in UV wind lines has been reported in colliding-wind binaries, where wind-wind interaction regions can lead to orbital phase-dependent features (Stevens, 1993). The applicability of this interpretation to MWC 656 is uncertain given the weak wind of the Be star. Alternatively, the interaction between the wind of the stripped star and the disc and / or the wind of the Be star could generate velocity components decoupled from orbital motion and could also account for the observed X-ray emission of the system (Langer et al., 2020a).
@@
Additional evidence for wind-disc interaction comes from the transient appearance of metastable He i∗ absorption lines, which are known tracers of outer-disc irradiation. In four ARCES spectra near quadrature (–0.44), we detect sharp absorption in He i , consistent with heating of the Be star’s disc by the hot companion. Further observations are needed to test this scenario; in particular, a UV spectrum obtained near inferior conjunction – when the companion passes in front of the Be star – would probe the hemisphere of the stripped star facing away from the Be star’s radiation, where wind-disc interaction is expected to be weaker.
@@
Another possible signature of interaction is the deep, narrow, near-rest-frame absorption component observed in N v (Fig. 4). Its depth is too large to be explained by a stripped companion contributing only 10–15% of the total flux, instead requiring a substantial fraction of the Be-star light to be absorbed by low-velocity, ionised material along the line of sight. This material may be associated with the Be star that is irradiated by the hot companion. An origin intrinsic to the Be star itself appears unlikely, as N v is not expected for a mid-type Be star; this is also illustrated by comparison with other Be+subdwarf systems (Fig. 8), none of which show similarly strong, narrow N v absorption, with only a partial exception in the earlier-type system 60 Cyg. Moreover, the absorption does not follow the Be star’s orbital motion. While the physical origin of this absorption component remains uncertain, it does not affect our main conclusion that a hot, stripped, luminous companion provides the most plausible explanation for the underlying N v P-Cygni profile, including its emission and blue-shifted absorption trough.
@@ 9 Summary and conclusions
Using new high-resolution optical spectra, we revised the orbital solution of the system. Our updated binary parameters yield minimum dynamical masses of M⊙ for the Be star and for its companion. A detailed spectroscopic analysis of the Be star allows us to constrain its effective temperature, luminosity, and mass more robustly, resulting in an estimated spectroscopic mass of . Together with the dynamical constraints, this implies a companion mass of approximately .
@@
To further investigate the nature of the companion, we obtained multiepoch HST/STIS far-UV spectra of MWC 656. These reveal high-ionisation wind features, such as N v lines with P Cygni-like profiles, which cannot originate in the Be star. The line strengths vary with orbital phase, increasing near quadrature and becoming weaker near conjunction when the companion is behind the Be star. This behaviour suggests partial obscuration of the line-forming region and supports an origin in the hot wind of a luminous companion.
We considered alternative scenarios in which the companion is a compact object accreting from the Be star’s wind. However, we find that the level of X-ray emission required to power the observed wind features is inconsistent with the observed X-ray luminosity of the system, and such models would predict additional spectral features, such as strong C iv emission, that are not observed.
@@
Although the nature of RV variability in the wind lines and origin of the deep, narrow absorption component in N v remain uncertain, the presence of a hot stripped companion provides a coherent explanation for most of the UV features. This interpretation not only rules out the BH scenario but highlights MWC 656 as a valuable system for studying the population of intermediate-mass stripped stars. Depending on the precise mass of the Be star, the system may eventually evolve into a double white dwarf or neutron star binary, making it a compelling target for future population synthesis and evolutionary studies.
@@
The newly calculated PoWR model spectra for the Be star and the companion will be provided as supplementary material on Zenodo. The HST UV spectra used in this work are publicly accessible and can be obtained from MAST. All other spectroscopic data will be shared upon reasonable request to the authors.
@@ Acknowledgements.
- A Be star-black hole binary with a wide orbit from LAMOST time-domain survey. arXiv e-prints, pp. arXiv:2505.23151. External Links: Document, 2505.23151, ADS entry Cited by: §1.
@@
- The Stripped-Star Ultraviolet Magellanic Cloud Survey (SUMS): The UV Photometric Catalog and Stripped Star Candidate Selection. arXiv e-prints, pp. arXiv:2505.18632. External Links: Document, 2505.18632, ADS entry Cited by: §8.1.
@@
- Minimum and maximum mass-luminosity relations for stripped stars. arXiv e-prints, pp. arXiv:2508.14161. External Links: Document, 2508.14161, ADS entry Cited by: Appendix F.
@@
- Discovery of a new transitional type of evolved massive stars with hard ionizing flux. arXiv e-prints, pp. arXiv:2508.18410. External Links: Document, 2508.18410, ADS entry Cited by: §6.
@@
- Populations of evolved massive binary stars in the Small Magellanic Cloud I: Predictions from detailed evolution models. arXiv e-prints, pp. arXiv:2503.23876. External Links: Document, 2503.23876, ADS entry Cited by: §1.
@@ Appendix A Stellar abundances
The abundances adopted in the composite spectral analysis are given in Table 3. For the Be star, C, N, and O mass fractions were adjusted to match diagnostic lines in the optical and UV spectra. Here, we focused in particular on N ii –––, O ii , O iii –, C ii and C iii –. For the stripped companion, which lacks spectroscopic constraints, we adopted fixed values guided by analyses of known stripped stars. For other elements, we assumed solar abundances as derived by Asplund et al. (2009) and implemented in Hainich et al. (2019).
@@ Appendix B Spectral disentangling in the UV
We performed spectral disentangling of the UV spectra using the same approach as for the optical spectra (see Sect. 3.3), aiming to separate the contributions of the Be star and its companion. Here, we focused on the wavelength range 1635–1650 Å, where there is clearly visible phase-dependent variability (see Fig. 4).
@@ Appendix C Stellar parameters of the Be star
Figure 12 shows a close-up comparison between the observed optical spectrum of MWC 656 and the best-fit model for the Be star. Highlighted are several diagnostic lines used to infer the stellar parameters, including the effective temperature, surface gravity, and projected rotational velocity.
@@ Appendix D Observation overview
| ARCES | 57643.23 | 150.8 | 0.34 | 45.67 | 4.10 | 5.77 | |
| ARCES | 58000.29 | 103.4 | 0.39 | 73.35 | 5.56 | 5.45 | |
| ARCES | 59540.19 | 173.7 | 0.48 | 57.35 | 4.04 | 4.43 | |
| ARCES | 59545.15 | 148.7 | 0.54 | 53.17 | 3.88 | 4.20 | |
| ARCES | 58085.15 | 153.2 | 0.83 | 25.45 | 4.10 | 3.90 | |
| ARCES | 59007.38 | 128.5 | 0.45 | 42.72 | 3.83 | 4.04 | |
| ARCES | 59015.41 | 136.3 | 0.59 | 39.54 | 4.02 | 3.36 | |
| ARCES | 59186.22 | 128.9 | 0.48 | 47.72 | 4.12 | 3.89 | |
| ARCES | 59214.10 | 105.5 | 0.95 | 29.37 | 7.07 | 1.68 | |
| ARCES | 59216.09 | 202.2 | 0.99 | 39.28 | 4.06 | 1.27 | |
| ARCES | 59536.22 | 167.3 | 0.41 | 73.68 | 4.08 | 5.41 | |
| ARCES | 59538.04 | 207.4 | 0.44 | 58.02 | 3.99 | 3.94 | |
| ESPaDOnS | 57201.55 | 373.2 | 0.63 | 43.79 | 4.55 | 2.24 | |
| ESPaDOnS | 57208.57 | 379.9 | 0.98 | 43.76 | 3.37 | 0.94 | |
| ESPaDOnS | 57227.57 | 217.8 | 0.30 | 67.50 | 3.30 | 3.32 | |
| HIRES | 60189.36 | 198.2 | 0.47 | 29.85 | 3.28 | 2.41 | |
| HIRES | 60213.36 | 178.7 | 0.88 | 17.93 | 4.11 | 2.25 | |
| HIRES | 60271.19 | 176.8 | 0.86 | 22.94 | 3.08 | 2.09 | |
| HIRES | 60274.30 | 186.3 | 0.91 | 22.16 | 3.85 | 1.81 | |
| HIRES | 60275.37 | 144.9 | 0.93 | 27.81 | 3.40 | 1.63 | |
| TRES | 59857.26 | 39.3 | 0.85 | 52.67 | 12.03 | 5.88 | |
| TRES | 59864.28 | 50.0 | 0.96 | 28.15 | 7.66 | 3.33 | |
| TRES | 59870.22 | 59.5 | 0.07 | 53.98 | 9.05 | 5.21 | |
| TRES | 59879.29 | 70.9 | 0.22 | 65.78 | 8.39 | 3.61 | |
| TRES | 59885.26 | 66.7 | 0.32 | 56.69 | 8.03 | 4.67 | |
| TRES | 59891.19 | 50.3 | 0.42 | 72.19 | 8.85 | 5.67 | |
| TRES | 59900.26 | 25.0 | 0.57 | 52.38 | 12.82 | 8.48 | |
| TRES | 59912.17 | 29.8 | 0.78 | 27.81 | 13.48 | 8.19 | |
| TRES | 59922.22 | 42.2 | 0.95 | 50.91 | 10.57 | 6.15 | |
@@
| HST/STIS | 59992.7 | 40 | 0.14 | — | |||
| HST/STIS | 59920.6 | 46 | 0.92 | — | |||
| HST/STIS | 59947.7 | 44 | 0.38 | — | |||
| HST/STIS | 60347.3 | 45 | 0.15 | — | |||
| HST/COS | 59992.7 | 39 | 0.14 | — | |||
| HST/STIS (FUV) | 59934.0 | 10 | 0.15 | — | |||
@@ Appendix F Circumsecondary accretion disc
The geometry of the system supports the possibility of mass transfer. From the H equivalent width and using the algorithm of Grundstrom and Gies (2006), we estimated the Be star’s disc half-width at half-maximum emission radius to be , comparable to its Roche-lobe radius (). This suggests that the outer Be disc can extend far enough to lose gas toward the secondary. Numerical simulations of Be + stripped-star systems indicate that such transfer can form Keplerian accretion discs around the companion (e.g. Panoglou et al. 2016; Rubio et al. 2025), and predict X-ray luminosities consistent with those observed in MWC 656 (Rast et al. 2025). Assuming the He ii line-peak separation traces Keplerian motion, we infer an outer radius of (, where is the orbital semi-major axis), similar to the Roche-lobe radius of the companion. Such a disc around the companion could contribute to the UV flux and host very hot plasma ( K; Peters and Polidan 1984), potentially obscuring part of the companion’s flux and producing redshifted N v absorption, as seen in HR 2142 (Peters et al. 2016). However, the observed He ii emission lines are narrower than expected for a Keplerian disc around a high-gravity () star – the inferred inner-disc rotation speed would exceed the measured width by approximately a factor of two. Moreover, a circumsecondary disc cannot reproduce the P Cygni-like N v profile, arguing against the disc as the dominant source of the optical He ii emission.

### v2

@@ 8.2 Caveats of the composite model
A notable feature of the UV data is an apparent Doppler shift in the high-ionisation wind lines. In three of the four epochs, the lines show no significant shift, but at the red edge of the absorption component, clearly seen, for example, in C iv (Fig. 4), appears redshifted by 80 km s-1. The measurement of RVs in wind lines is inherently difficult due to their complex and variable profiles, and the situation could be further complicated here by interstellar absorption that overlaps with the wind features. Nevertheless, the observed shift cannot easily be explained by the orbital motion of either star: The amplitude is too small for the companion, the sign is opposite to the expected motion at this phase, and the shift is too large to be attributed to the Be star. This suggests that the high-ionisation line-forming region does not simply trace the orbital motion of either component. Redshifted absorption in UV wind lines has been reported in colliding-wind binaries, where wind-wind interaction regions can lead to orbital phase-dependent features (Stevens, 1993). The applicability of this interpretation to MWC 656 is uncertain given the weak wind of the Be star. Alternatively, the interaction between the wind of the stripped star and the disc and/or the wind of the Be star could generate velocity components decoupled from orbital motion and could also account for the observed X-ray emission of the system (Langer et al., 2020a).
@@
Additional evidence of wind–disc interaction comes from the transient appearance of metastable He i∗ absorption lines, which are known tracers of outer-disc irradiation. In four ARCES spectra near quadrature (–0.46), we detect sharp absorption in He i , consistent with heating of the Be star’s disc by the hot companion. Further observations are needed to test this scenario; in particular, a UV spectrum obtained near inferior conjunction – when the companion passes in front of the Be star – would probe the hemisphere of the stripped star facing away from the Be star’s radiation, where wind-disc interaction is expected to be weaker.
@@
Another possible signature of interaction is the deep, narrow, near-rest frame absorption component observed in N v (Fig. 4). Its depth is too large to be explained by a stripped companion contributing only 10–15% of the total flux, instead requiring a substantial fraction of the Be-star light to be absorbed by low-velocity, ionised material along the line of sight. This material may be associated with the Be star that is irradiated by the hot companion. An origin intrinsic to the Be star itself appears unlikely, as N v is not expected for a mid-type Be star; this is also illustrated by comparison with other Be+subdwarf systems (Fig. 8), none of which show similarly strong, narrow N v absorption, with only a partial exception in the earlier-type system 60 Cyg. Moreover, the absorption does not follow the orbital motion of the Be star. Although the physical origin of this absorption component remains uncertain, it does not affect our main conclusion that a hot, stripped, luminous companion provides the most plausible explanation for the underlying N v P-Cygni profile, including its emission and blueshifted absorption trough.
@@ 9 Summary and conclusions
Using new high-resolution optical spectra, we revised the orbital solution of the system. Our updated binary parameters yield minimum dynamical masses of M⊙ for the Be star and for its companion. A detailed spectroscopic analysis of the Be star allows us to constrain its effective temperature, luminosity, and mass more robustly, resulting in an estimated spectroscopic mass of . Together with dynamical constraints, this implies a companion mass of approximately .
@@
To further investigate the nature of the companion, we obtained multi-epoch HST/STIS FUV spectra of MWC 656. These reveal high-ionisation wind features, such as N v lines with P Cygni-like profiles, which cannot originate in the Be star. The line strengths vary with orbital phase, increasing near quadrature and becoming weaker near conjunction when the companion is behind the Be star. This behaviour suggests partial obscuration of the line-forming region and supports an origin in the hot wind of a luminous companion.
We considered alternative scenarios in which the companion is a compact object accreting from the Be star’s wind. However, we find that the level of X-ray emission required to power the observed wind features is inconsistent with the observed X-ray luminosity of the system, and such models would predict additional spectral features, such as strong C iv emission, which are not observed.
@@
Although the nature of RV variability in the wind lines and the origin of the deep and narrow absorption component in N v remain uncertain, the presence of a hot stripped companion provides a coherent explanation for most of the UV features. This interpretation not only rules out the BH scenario, but highlights MWC 656 as a valuable system for studying the population of intermediate-mass stripped stars. Depending on the precise mass of the Be star, the system may eventually evolve into a double white dwarf or neutron star binary, making it a compelling target for future population synthesis and evolutionary studies.
@@
The newly calculated PoWR model spectra for the Be star and the companion will be provided as supplementary material on Zenodo. The HST UV spectra used in this work are publicly accessible and can be obtained from MAST. Optical TRES and HIRES spectra are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
- The Stripped-Star Ultraviolet Magellanic Cloud Survey (SUMS): The UV Photometric Catalog and Stripped Star Candidate Selection. arXiv e-prints, submitted to ApJ, pp. arXiv:2505.18632. External Links: Document, 2505.18632, ADS entry Cited by: §8.1.
@@
- Minimum and maximum mass-luminosity relations for stripped stars. A&A 703, pp. A40. External Links: Document, 2508.14161, ADS entry Cited by: Appendix F.
@@
- Discovery of a new transitional type of evolved massive stars with hard ionizing flux. Nature Astronomy, Advanced Online Publication. External Links: Document, ADS entry Cited by: §6.
@@
- Populations of evolved massive binary stars in the Small Magellanic Cloud: I. Predictions from detailed evolution models. A&A 704, pp. A218. External Links: Document, 2503.23876, ADS entry Cited by: §1.
@@ Appendix A Stellar abundances
The abundances adopted in the composite spectral analysis are given in Table 3. For the Be star, C, N, and O mass fractions were adjusted to match diagnostic lines in the optical and UV spectra. Here, we focussed in particular on N ii –––, O ii , O iii –, C ii and C iii –. For the stripped companion, which lacks spectroscopic constraints, we adopted fixed values guided by analyses of known stripped stars. For other elements, we assumed solar abundances as derived by Asplund et al. (2009) and implemented in Hainich et al. (2019).
@@ Appendix B Spectral disentangling in the UV
We performed spectral disentangling of the UV spectra using the same approach as for the optical spectra (see Sect. 3.3), aiming to separate the contributions of the Be star and its companion. Here, we focussed on the wavelength range 1635–1650 Å, where there is clearly visible phase-dependent variability (see Fig. 4).
@@ Appendix C Stellar parameters of the Be star
Figure 12 shows a close-up comparison between the observed optical spectrum of MWC 656 and the best-fit model for the Be star. Highlighted are several diagnostic lines that were used to infer the stellar parameters, including effective temperature, surface gravity, and projected rotational velocity.
@@ Appendix D Observation overview
| ARCES | 57643.23 | 150.8 | 0.36 | 3.87 | 5.77 | ||
| ARCES | 58000.29 | 103.4 | 0.41 | 4.17 | 5.45 | ||
| ARCES | 59540.19 | 173.7 | 0.49 | 4.04 | 4.43 | ||
| ARCES | 59545.15 | 148.7 | 0.56 | 3.72 | 4.20 | ||
| ARCES | 58085.15 | 153.2 | 0.85 | 3.90 | 3.90 | ||
| ARCES | 59007.38 | 128.5 | 0.47 | 3.93 | 4.04 | ||
| ARCES | 59015.41 | 136.3 | 0.60 | 4.18 | 3.36 | ||
| ARCES | 59186.22 | 128.9 | 0.50 | 4.42 | 3.89 | ||
| ARCES | 59214.10 | 105.5 | 0.97 | 6.57 | 1.68 | ||
| ARCES | 59216.09 | 202.2 | 0.00 | 3.81 | 1.27 | ||
| ARCES | 59536.22 | 167.3 | 0.43 | 3.34 | 5.41 | ||
| ARCES | 59538.04 | 207.4 | 0.46 | 3.50 | 3.94 | ||
| ESPaDOnS | 57201.55 | 373.2 | 0.88 | 3.04 | 2.24 | ||
| ESPaDOnS | 57208.57 | 379.9 | 1.00 | 3.01 | 0.94 | ||
| ESPaDOnS | 57227.57 | 217.8 | 0.32 | 2.80 | 3.32 | ||
| HIRES | 60189.36 | 198.2 | 0.49 | 3.68 | 2.41 | ||
| HIRES | 60213.36 | 178.7 | 0.90 | 4.11 | 2.25 | ||
| HIRES | 60271.19 | 176.8 | 0.88 | 3.65 | 2.09 | ||
| HIRES | 60274.30 | 186.3 | 0.93 | 3.61 | 1.81 | ||
| HIRES | 60275.37 | 144.9 | 0.95 | 3.36 | 1.63 | ||
| TRES | 59857.26 | 39.3 | 0.86 | 9.98 | 5.88 | ||
| TRES | 59864.28 | 50.0 | 0.98 | 7.80 | 3.33 | ||
| TRES | 59870.22 | 59.5 | 0.08 | 8.18 | 5.21 | ||
| TRES | 59879.29 | 70.9 | 0.24 | 7.31 | 4.65 | ||
| TRES | 59885.26 | 66.7 | 0.34 | 6.68 | 4.67 | ||
| TRES | 59891.19 | 50.3 | 0.44 | 7.89 | 5.67 | ||
| TRES | 59900.26 | 25.0 | 0.59 | 11.85 | 8.48 | ||
| TRES | 59912.17 | 29.8 | 0.79 | 11.65 | 8.19 | ||
| TRES | 59922.22 | 42.2 | 0.96 | 10.59 | 6.15 | ||
@@
| HST/STIS | 59992.7 | 40 | 0.16 | — | |||
| HST/STIS | 59920.6 | 46 | 0.94 | — | |||
| HST/STIS | 59947.7 | 44 | 0.40 | — | |||
| HST/STIS | 60347.3 | 45 | 0.17 | — | |||
| HST/COS | 59992.7 | 39 | 0.16 | — | |||
| HST/STIS (FUV) | 59934.0 | 10 | 0.17 | — | |||
@@ Appendix F Circumsecondary accretion disc
The geometry of the system supports the possibility of mass transfer. From the H equivalent width and using the algorithm of Grundstrom and Gies (2006), we estimated the Be star’s disc half-width at half-maximum emission radius to be , comparable to its Roche-lobe radius (). This suggests that the outer Be disc can extend far enough to lose gas towards the secondary. Numerical simulations of Be + stripped-star systems indicate that such a transfer can form Keplerian accretion discs around the companion (e.g. Panoglou et al. 2016; Rubio et al. 2025), and predict X-ray luminosities consistent with those observed in MWC 656 (Rast et al. 2025). Assuming that the He ii line-peak separation traces Keplerian motion, we infer an outer radius of (, where is the orbital semi-major axis), similar to the Roche-lobe radius of the companion. Such a disc around the companion could contribute to the UV flux and host very hot plasma ( K; Peters and Polidan 1984), potentially obscuring part of the companion’s flux and producing redshifted N v absorption, as seen in HR 2142 (Peters et al. 2016). However, the observed He ii emission lines are narrower than expected for a Keplerian disc around a high-gravity () star – the inferred inner-disc rotation speed would exceed the measured width by approximately a factor of two. Moreover, a circumsecondary disc cannot reproduce the P Cygni-like N v profile, arguing against the disc as the dominant source of the optical He ii emission.